OpenStudy (anonymous):

Find cos θ if sin θ = -(5/13) and tan θ > 0

OpenStudy (jdoe0001):

$$\bf sin(\theta)= -\cfrac{5}{13} \implies \cfrac{opposite}{hypotenuse} \implies \frac{b}{c}\\ c^2=a^2+b^2 \implies b= \pm\sqrt{c^2-a^2} \implies b= \pm\sqrt{144} \implies b=12\\ \text{now } tan(\theta) > 0 \text{ meaning is positive} $$ so for \( tan(\theta) to be positive it has to have either sine or cosine negative and the other positve as we can see above, sine is already negative \(\bf in(\theta)= -\cfrac{5}{13} so to get a positive tangent, cosine has to also be negative, thus b = -12

OpenStudy (jdoe0001):

hmm, lemme redo that :/

OpenStudy (jdoe0001):

$$\bf sin(\theta)= -\cfrac{5}{13} \implies \cfrac{opposite}{hypotenuse} \implies \frac{b}{c}\\ c^2=a^2+b^2 \implies b= \pm\sqrt{c^2-a^2} \implies b= \pm\sqrt{144} \implies b=12\\ \text{now } tan(\theta) > 0 \text{ meaning is positive} $$ so for \( tan(\theta)\) to be positive it has to have either sine or cosine negative and the other positve as we can see above, sine is already negative, \(\bf sin(\theta)= -\cfrac{5}{13}\) so to get a positive tangent, cosine has to also be negative, thus b = -12

OpenStudy (anonymous):

Haha i understood what you were saying :P no need to redo. and that's what i thought. so it would be -12/13 correct?

OpenStudy (jdoe0001):

wait a second, I'm after "a", NOT b component

OpenStudy (anonymous):

oh :P

OpenStudy (jdoe0001):

well, the same value will come out anyway :)

OpenStudy (jdoe0001):

\(\bf sin(\theta)= -\cfrac{5}{13} \implies \cfrac{opposite}{hypotenuse} \implies \frac{b}{c}\\ c^2=a^2+b^2 \implies a= \pm\sqrt{c^2-b^2} \implies a= \pm\sqrt{144} \implies a=12\)

OpenStudy (anonymous):

So it should be -12/13 correct?

OpenStudy (jdoe0001):

and it will be -12 because the tangent is positive, so the adjacent side will also be negative

OpenStudy (jdoe0001):

yes, is correct

OpenStudy (anonymous):

 thats what i get though?